OpenStudy (calculusxy):

The sum of two consecutive add numbers is 6 less than twice 5 squared, what are the numbers? @jdoe0001

OpenStudy (jdoe0001):

if a number is say "a" the consecutive number can be 1 number away from it, so that can be a-1 or a+1 let's use a+1 so \(\bf a + (a+1) = \) 5squared, twice => \(\bf 2 \times 5^2\) 6 less than that, = >\(\bf 2 \times 5^2 -6\) so \(\bf a + (a+1) = 2 \times 5^2 -6\implies 2a + 1 = 50-6\)